OpenStudy (rainbow_rocks03):

A spinner has 10 equal sections labeled 1 through 10. A bag contains five letter tiles labeled A, A, B, B, and B. What is the probability of spinning a multiple of 3 on the spinner and drawing a letter tile labeled B from the bag?

OpenStudy (anonymous):

1/10

OpenStudy (kendricklamar2014):

Probability of spinning a multiple of three is:\[\frac{ 3 }{ 10 }\] Probability of drawing a B is:\[\frac{ 3 }{ 5 }\]

OpenStudy (geekfromthefutur):

What do u need help with?

OpenStudy (rainbow_rocks03):

these are the answer choices

OpenStudy (kendricklamar2014):

Probability of spinning a multiple of 3 and drawing a B is:\[\frac{ 3 }{ 10 }\times \frac{ 3 }{ 5 }\]

OpenStudy (kendricklamar2014):

\[= \frac{ 9 }{ 50 }\]

OpenStudy (rainbow_rocks03):

a.2/5 b.6/25 c.9/50 d.3/20

OpenStudy (rainbow_rocks03):

thx

OpenStudy (kendricklamar2014):

You're Welcome